OpenStudy (anonymous):

In a college biology class, only half of the class earned a passing grade. Out of all the students, 45% studied and passed. What is the probability of someone having studied, given that they passed A) 1% B) 55% C) 90% D) 45%

OpenStudy (anonymous):

.45/.5 = 0.9 or 90%

OpenStudy (anonymous):

45% studied and passed. only half of the class earned a passing grade. (.5) P(studied and pass test) / P(passed) = 0.45/ 0.5 = 0.9 or 90%